OpenStudy (zepp):

I'm trying some summation this afternoon, I was trying to add up positive odd numbers. So 1+3+5+7+9+11+13... etc. etc. Odd number could be written algebraically as 2x+1, x being any positive natural number. \[\huge \sum_{x=0}^{n} (2x+1)\] And this is what I got: \[\huge \sum_{x=0}^{n} (2x+1)=(n-1)^2, n \in \Natural\]

OpenStudy (zepp):

Now I'm trying \[\huge \sum_{x=0}^{n}\frac{1}{2x+1}\]

OpenStudy (zepp):

but I can't find how to write that in with n as variable :|

OpenStudy (blockcolder):

So in short, you're trying to find a simpler expression for \[\sum_{x=0}^n \frac{1}{2x+1}\] in terms of n but you can't find one, right?

OpenStudy (zepp):

Yup :)

OpenStudy (anonymous):

Not possible without using some advance function.

OpenStudy (anonymous):

And your earlier answer is incorrect, \[ \sum_{x=0}^{n} (2x+1)=(n\color{red}{+}1)^2, n \in \mathbb{N} \]

OpenStudy (zepp):

Oh, I should check my work again D:

OpenStudy (anonymous):

To be more pedantic: \[ \sum_{x=0}^{n} (2x+1)=(n\color{red}{+}1)^2, n \in \mathbb{N_0} \]

OpenStudy (blockcolder):

Let \(H_n\) be the nth harmonic number. http://en.wikipedia.org/wiki/Harmonic_number \[\begin{align}1+\frac{1}{3}+\frac{1}{5}+\cdots+\frac{1}{2n+1}&=H_{2n+2}-\left(\frac{1}{2}+\frac{1}{4}+\cdots+\frac{1}{2n}\right)\\ &=H_{2n+2}-\frac{H_n}{2} \end{align}\] This was the first thing I thought of.